# Surveying the Landscape of Image Captioning Evaluation: A Comprehensive Taxonomy, Trends and Metrics Analysis

### v2

@@ Abstract
The task of image captioning has recently been gaining popularity, and with it the complex task of evaluating the quality of image captioning models. In this work, we present the first survey and taxonomy of over 70 different image captioning metrics and their usage in hundreds of papers, specifically designed to help users select the most suitable metric for their needs. We find that despite the diversity of proposed metrics, the vast majority of studies rely on only five popular metrics, which we show to be weakly correlated with human ratings. We hypothesize that combining a diverse set of metrics can enhance correlation with human ratings. As an initial step, we demonstrate that a linear regression-based ensemble method, which we call EnsembEval, trained on one human ratings dataset, achieves improved correlation across five additional datasets, showing there is a lot of room for improvement by leveraging a diverse set of metrics. [^1] Our code and data are available on
@@ 1 Introduction
Figure 1: An image with various valid captions that have no overlap in content words, exemplifying some of the challenges in evaluating image captioning.
@@ 2 Task Definition
Early research framed image captioning as a means to assess vision-and-language models based on image-caption semantics rather than practical application Hodosh et al. (2013). Recently the task is commonly defined as describing the visual content of an image or scene in natural language Stefanini et al. (2022); Nallapaneni and Konakanchi (2023); González-Chávez et al. (2024). Yet, this definition does not specify what visual content should be described or the intended nature of the description. These aspects depend on both the purpose and the context of the description, which we now discuss in detail.
Descriptions serve different purposes, each with distinct requirements. For example, captions designed for accessibility should replace the image, while captions in news articles typically augment or link the content of the image and article Kreiss et al. (2022). In practice, although often implicit, most previous studies focus on image replacement, i.e., generating descriptions intended to serve as substitutes for the image, as defined by Kreiss et al. (2022).
The second key factor is the context in which a caption is generated. For instance, the top caption in Figure 1 may suit Lego’s Instagram account, while the center caption might be more appropriate for a travel guide.
While our survey covers all works on captioning evaluation – irrespective of purpose or context – our experiments on metric performance and ensembling (Section 6) focus on the replacement task, specifically .
@@ 4.1.1 Relation with Ground-Truth
Phrase-level semantic similarity: Others compare the semantics of sentence elements, most commonly using phrase embeddings, either with context (BERTScoreZhang et al. 2019, BERTScore++Yi et al. 2020) or without (WMDKusner et al. 2015, NWCornia et al. 2019a, SSDTakmaz et al. 2020, SemSimNag Chowdhury et al. 2021, Fuzzy noun/verb overlapChan et al. 2023a). VILBERTScore Lee et al. (2020) enriches phrase embeddings by injecting visual information. The SPICE family (SPICEAnderson et al. 2016, SPICE-UWang et al. 2020d) compares the components of scene graphs of the candidate and reference captions. ALOHa Petryk et al. (2024) prompts Large Language Models (LLMs) to identify object phrases in both candidate and references and computes the similarity of these phrases’ embeddings.
Sentence-level similarity: Some metrics compare semantics at the sentence level. MPNetScore Black et al. (2024a) compares sentence level embeddings. Others (CLAIRChan et al. 2023b, GPT4V evaluationGe et al. 2024a) prompt LLMs to compare the candidate and reference caption. Nubia Kane et al. (2020) explicitly trains a model to evaluate sentence similarity.
@@
Sentence-image similarity: Several recent works embeded sentences and images in a joint visual-textual space and computed the similarity of candidate-image embeddings (Semantic ScoreDognin et al. 2019, CLIPScoreHessel et al. 2021, PR-MCSKim et al. 2023, PACScoreSarto et al. 2023, BLIP2ScoreZeng et al. 2024a). CLIPImageScore Ge et al. (2024a) use text-to-image methods to generate an image based on the candidate and computes the embedding similarity between this generated image and the original image. UMIC Lee et al. (2021b) explicitly trains models to predict similarity of candidate caption and image.
@@ 4.1.3 Candidate Diversity
Lexical diversity: While some works use lexical similarity methods such as BLEU (mBLEUShetty et al. 2017, self-BLEUZhu et al. 2018) and CIDEr (Self-CIDErWang and Chan 2019, CIDErBtwWang et al. 2020a) on pairs of captions from $S$, others specifically designed methods to measure lexical diversity in $S$ (VocabShetty et al. 2017, Div-$n$Shetty et al. 2017, diversity edit distanceDai et al. 2018a, TdivLiu et al. 2019a, DistLiu et al. 2019a).
@@ 6.3 EnsembEval: An Ensemble of Evaluation Methods
Table 5: Accuracy in pairwise comparison. In Pascal50S we report mean and standard deviation across five random instances of tie-breaking and reference selection. HC: both captions are human-generated for the image, HI: both are human-generated but only one corresponds to the image, HM: one caption is human-generated while the other is machine-generated, MM: both captions are machine-generated. In each dataset, best scoring metric is in bold.
@@ 8 Conclusion
We conducted a comprehensive survey of image captioning evaluation methods, yielding two key outcomes: a taxonomy of captioning metrics, including overlooked high-performing ones (e.g., BLIP2Score) and an analysis revealing that most papers rely on five metrics with weak human correlation. We further showed that an ensemble of metrics improves correlation with human ratings, highlighting the potential for further advancements in this approach.
In addition to the theoretical contribution, we anticipate two practical uses for EnsembEval. First, it is readily available131313github.com/uriberger/caption_evaluation as a reliable metric with strong correlation to human ratings. Second, users can extend the ensemble by incorporating additional metrics or adjusting the weights.
@@ Appendix B Metric Usage Records
Vision Ge et al. (2024b): RefSenSim, SenImSim, Huang et al. (2024): LexSim, RefPhSim, Zeng et al. (2024b): LexSim, RefPhSim, SenImSim, Li et al. (2024a): LexSim, RefPhSim
Machine Learning Ma et al. (2024b): LexSim, RefPhSim, Black et al. (2024b): LexSim, RefSenSim, Liu et al. (2024): LexSim, RefPhSim, Ma et al. (2024a): LexSim, Qi et al. (2024): LexSim, Qiu et al. (2024): LexSim, RefPhSim, Wang et al. (2024): LexSim, RefPhSim, Fu et al. (2024): LexSim, RefPhSim
@@
Vision Wu et al. (2023a): LexSim, Li et al. (2023b): Flu, Div, Kornblith et al. (2023): LexSim, SenImSim, Ret, MulSim, Hu et al. (2023b): LexSim, RefPhSim, Tu et al. (2023): LexSim, RefPhSim, Fei et al. (2023a): LexSim, RefPhSim, Fan et al. (2023): LexSim, RefPhSim, Kang et al. (2023): LexSim, RefPhSim, SenImSim, Ret, Barraco et al. (2023): LexSim, RefPhSim, SenImSim, MulSim, Hu et al. (2023b): LexSim, RefPhSim, Chen et al. (2023b): LexSim, Dessì et al. (2023): LexSim, RefPhSim, Vo et al. (2023): LexSim, RefPhSim, SenImSim, Ret, MulSim, Luo et al. (2023a): LexSim, RefPhSim, Zeng et al. (2023): LexSim, RefPhSim, SenImSim, MulSim, Div, Kuo and Kira (2023): LexSim, RefPhSim, Chen et al. (2023a): LexSim, Ramos et al. (2023c): LexSim, RefPhSim, Hirota et al. (2023): LexSim, RefPhSim, SenImSim, Bias, Ren et al. (2023): LexSim, RefPhSim
@@
Machine Learning Del Chiaro et al. (2020): LexSim, Mahajan and Roth (2020): LexSim, RefPhSim, Div, Zhang et al. (2020): LexSim, RefPhSim, Zhou et al. (2020a): LexSim, RefPhSim, Zhao et al. (2020): LexSim, Flu, Wang et al. (2020c): LexSim, RefPhSim, Liu et al. (2020): LexSim, Hou et al. (2020): LexSim, RefPhSim, Cao et al. (2020): LexSim, RefPhSim

### v3

@@ Abstract
The task of image captioning has recently been gaining popularity, and with it the complex task of evaluating the quality of image captioning models. In this work, we present the first survey and taxonomy of over 70 different image captioning metrics and their usage in hundreds of papers, specifically designed to help users select the most suitable metric for their needs. We find that despite the diversity of proposed metrics, the vast majority of studies rely on only five popular metrics, which we show to be weakly correlated with human ratings. We hypothesize that combining a diverse set of metrics can enhance correlation with human ratings. As an initial step, we demonstrate that a linear regression-based ensemble method, which we call EnsembEval, trained on one human ratings dataset, achieves improved correlation across five additional datasets, showing there is a lot of room for improvement by leveraging a diverse set of metrics. [^1] Our code and data are available at
@@ 1 Introduction
Figure 1: An image with various valid captions that have no overlap in content words, exemplifying some of the challenges in evaluating image captioning (taken from Crossmodal-3600; Thapliyal et al., 2022).
@@ 2 Task Definition
Early research framed image captioning as a means to assess vision-and-language models based on image-caption semantics, rather than practical applications Hodosh et al. (2013). More recently, the task has often been defined simply as describing the visual content of an image or scene in natural language Stefanini et al. (2022); Nallapaneni and Konakanchi (2023); González-Chávez et al. (2024). Yet, this definition does not specify what visual content should be described or the intended nature of the description. These aspects depend on both the purpose and the context of the description, which we now discuss in detail.
Descriptions serve different purposes, each with distinct requirements. For example, captions designed for accessibility should replace the image, while captions in news articles typically augment or link the content of the image and article Kreiss et al. (2022). In practice, most studies focus on image replacement, i.e., generating descriptions intended to serve as substitutes for the image, as defined by Kreiss et al. (2022).
The exact purpose of image captioning is often implicit in image captioning papers. Hodosh et al. (2013) discusses the task in detail, and adopts a number of fundamental theoretical distinctions as to the purposes of captioning, taken from library science (Shatford, 1986; Jaimes and Chang, 1999). Specifically, they argue that captions that focus on conceptual descriptions (rather than low-level “perceptual” descriptions) are the most relevant for image understanding, the most commonly tackled purpose in NLP. Within this category, an important sub-distinction is between specific captions, that identify entities by their names (e.g., Great Pyramid of Giza), and generic ones (e.g., that instead identify an object as a pyramid). Image understanding usually (albeit, not universally) targets the latter.
The second key factor is the context in which a caption is generated. For instance, the top caption in Figure 1 may suit Lego’s Instagram account, while the middle caption might be more appropriate for a travel guide.
While our survey covers all works on captioning evaluation — irrespective of purpose or context — our experiments on metric performance and ensembling (Section 6) focus on the replacement task, specifically.
@@ 4.1.1 Relation with Ground-Truth
Phrase-level semantic similarity: Others compare the semantics of sentence elements, most commonly using phrase embeddings, either with context (BERTScoreZhang et al. 2020a, BERTScore++Yi et al. 2020) or without (WMDKusner et al. 2015, NWCornia et al. 2019a, SSDTakmaz et al. 2020, SemSimNag Chowdhury et al. 2021, Fuzzy noun/verb overlapChan et al. 2023a). VILBERTScore Lee et al. (2020) enriches phrase embeddings by injecting visual information. The SPICE family (SPICEAnderson et al. 2016, SPICE-UWang et al. 2020d) compares the components of scene graphs of the candidate and reference captions. ALOHa Petryk et al. (2024) prompts Large Language Models (LLMs) to identify object phrases in both candidate and references and computes the similarity of these phrases’ embeddings.
Sentence-level similarity: Some metrics compare semantics at the sentence level. MPNetScore Black et al. (2024) compares sentence level embeddings. Others (CLAIRChan et al. 2023b, GPT4V evaluationGe et al. 2024a) prompt LLMs to compare the candidate and reference caption. Nubia Kane et al. (2020) explicitly trains a model to evaluate sentence similarity.
@@
Sentence-image similarity: Several recent works embeded sentences and images in a joint visual-textual space and computed the similarity of candidate-image embeddings (Semantic ScoreDognin et al. 2019, CLIPScoreHessel et al. 2021, PR-MCSKim et al. 2023, PACScoreSarto et al. 2023, BLIP2ScoreZeng et al. 2024). CLIPImageScore Ge et al. (2024a) use text-to-image methods to generate an image based on the candidate and computes the embedding similarity between this generated image and the original image. UMIC Lee et al. (2021b) explicitly trains models to predict similarity of candidate caption and image.
@@ 4.1.3 Candidate Diversity
Lexical diversity: While some works use lexical similarity methods such as BLEU (mBLEUShetty et al. 2017, self-BLEUZhu et al. 2018) and CIDEr (Self-CIDErWang and Chan 2019, CIDErBtwWang et al. 2020a) on pairs of captions from $S$, others specifically designed methods to measure lexical diversity in $S$ (VocabShetty et al. 2017, Div-nnShetty et al. 2017, diversity edit distanceDai et al. 2018a, TdivLiu et al. 2019a, DistLiu et al. 2019a).
@@ 6.3 EnsembEval: An Ensemble of Evaluation Methods
Table 5: Accuracy in pairwise comparison. In Pascal50S we report mean and standard deviation across five random instances of tie-breaking and reference selection. HC: both captions are human-generated for the image, HI: both are human-generated but only one corresponds to the image, HM: one caption is human-generated while the other is machine-generated, $M$: both captions are machine-generated. In each dataset, best scoring metric is in bold.
@@ 6.6 Why Does the Ensemble Improve Correlation?
7 Software Package
Our empirical analysis of the redundancy/complementarity of the different metrics, reveals that an ensemble of several methods may be considerably better correlated with human rankings of the captions. To this aim, we release a software package EnsembEval under the permissive MIT license.131313github.com/uriberger/caption_evaluation
The software package implements the ensemble metric presented in Section 6.3. It also straightforwardly supports the use of alternative sets of weights. The software package can easily be extended with additional metrics, either reference-based or reference-free. The package is therefore usable not only for replication of the results presented in this paper, but also as a generic tool for ensembling image captioning evaluation metrics.
@@ 9 Conclusion
We conducted a comprehensive survey of image captioning evaluation methods, yielding two key outcomes: a taxonomy of captioning metrics, including overlooked high-performing ones (e.g., BLIP2Score) and an analysis revealing that most papers rely on five metrics that have only a weak correlation with human ratings. We further showed that an ensemble of several existing metrics improves correlation with human ratings, and proposed a simple weighted ensemble method to this effect. We release a software package that implements this ensemble methods and facilitates future developments in image captioning metric ensembling methods.
@@ Appendix B Metric Usage Records
Vision Ge et al. (2024b): RefSenSim, SenImSim, Huang et al. (2024): LexSim, RefPhSim, Zeng et al. (2024): LexSim, RefPhSim, SenImSim, Li et al. (2024a): LexSim, RefPhSim
Machine Learning Ma et al. (2024b): LexSim, RefPhSim, Black et al. (2024): LexSim, RefSenSim, Liu et al. (2024): LexSim, RefPhSim, Ma et al. (2024a): LexSim, Qi et al. (2024): LexSim, Qiu et al. (2024): LexSim, RefPhSim, Wang et al. (2024): LexSim, RefPhSim, Fu et al. (2024): LexSim, RefPhSim
@@
Vision Wu et al. (2023a): LexSim, Li et al. (2023b): Flu, Div, Kornblith et al. (2023): LexSim, SenImSim, Ret, MulSim, Hu et al. (2023b): LexSim, RefPhSim, Tu et al. (2023): LexSim, RefPhSim, Fei et al. (2023a): LexSim, RefPhSim, Fan et al. (2023): LexSim, RefPhSim, Kang et al. (2023): LexSim, RefPhSim, SenImSim, Ret, Barraco et al. (2023): LexSim, RefPhSim, SenImSim, MulSim, Hu et al. (2023b): LexSim, RefPhSim, Chen et al. (2023b): LexSim, Dess`ı et al. (2023): LexSim, RefPhSim, Vo et al. (2023): LexSim, RefPhSim, SenImSim, Ret, MulSim, Luo et al. (2023a): LexSim, RefPhSim, Zeng et al. (2023): LexSim, RefPhSim, SenImSim, MulSim, Div, Kuo and Kira (2023): LexSim, RefPhSim, Chen et al. (2023a): LexSim, Ramos et al. (2023c): LexSim, RefPhSim, Hirota et al. (2023): LexSim, RefPhSim, SenImSim, Bias, Ren et al. (2023): LexSim, RefPhSim
@@
Machine Learning Del Chiaro et al. (2020): LexSim, Mahajan and Roth (2020): LexSim, RefPhSim, Div, Zhang et al. (2020b): LexSim, RefPhSim, Zhou et al. (2020a): LexSim, RefPhSim, Zhao et al. (2020): LexSim, Flu, Wang et al. (2020c): LexSim, RefPhSim, Liu et al. (2020): LexSim, Hou et al. (2020): LexSim, RefPhSim, Cao et al. (2020): LexSim, RefPhSim